# The Illusion of Readiness: Stress Testing Large Frontier Models on Multimodal Medical Benchmarks

### v2

@@ Abstract
Large frontier models like GPT-5 now achieve top scores on medical benchmarks. But our stress tests tell a different story. Leading systems often guess correctly even when key inputs like images are removed, flip answers under trivial prompt changes, and fabricate convincing yet flawed reasoning. These aren’t glitches; they expose how today’s benchmarks reward test-taking tricks over medical understanding. We evaluate six flagship models across six widely used benchmarks and find that high leaderboard scores hide brittleness and shortcut learning. Through clinician-guided rubric evaluation, we show that benchmarks vary widely in what they truly measure yet are treated interchangeably, masking failure modes. We caution that medical benchmark scores do not directly reflect real-world readiness. If we want AI to earn trust in healthcare, we must demand more than leaderboard wins and must hold systems accountable for robustness, sound reasoning, and alignment with real medical demands.
1 Introduction: The Illusion of Readiness in Health AI
Health AI has a credibility problem. Models like GPT-5 [1] can ace medical exams and top multimodal benchmarks yet still falter on simpler tasks, such as preserving their answers when answer choices are shuffled, or justifying their predictions with medically sound reasoning. This isn’t just an implementation lag. We’re measuring the wrong things.
Current health AI benchmarks reward test-taking strategies rather than robust medical cross-modal understanding [2]. A model might diagnose pneumonia not by interpreting radiologic features, but by learning that “productive cough + fever” statistically co-occurs with pneumonia in training data. This is shortcut learning, not medical understanding.
Figure 1: Stress tests reveal hidden fragilities in LMMs on multimodal medical tasks. a. Conceptual illustration: Benchmark scores suggest steady model improvement. Stress tests uncover hidden vulnerabilities—newer models may be equally or more brittle despite higher scores. b. While conventional benchmark scores (green line) show progress, robustness scores (red line) expose increasing brittleness under incomplete or adversarial inputs. c. Stress test heatmap disrupts the apparent leaderboard, revealing model-specific failure patterns. d. Illustrative examples: (1) correct answer despite missing image and shuffled choices (shortcut behavior); (2) fabricated rationale when asked to explain an answer without visual input.
Note: Panel d examples are generated from GPT-5 responses when prompted post hoc for reasoning. Response variability may occur across runs. To enforce direct answering, all prompts included the suffix: “Please only answer with the index and content of the option.” See Appendix E: Prompt Design.
When we looked deeper, the problems became clear:
  • Models succeed for the wrong reasons: On multimodal medical benchmarks [3, 4], leading models retained most of their original accuracy even when images were removed. For questions that explicitly require visual input, they still guess correctly, with cues as minimal as a single familiar distractor.
  • Brittle performance: Reordering answer choices, weakening distractors, or subtly changing the image causes large shifts in predictions, despite no change to the core medical question.
  • Fabricated reasoning: Models trained to “think step by step” [5] often paired confident rationales with incorrect logic producing medically sound explanations for wrong answers, or correct answers supported by fabricated reasoning.
These aren’t minor technical glitches. They reveal fundamental problems with how we evaluate and incentivize progress in health AI. Current benchmarks test pattern matching, not medical understanding. They reward consistency on test formats rather than robustness under real medical conditions.
Real-world medical decisions are made under uncertainty, incomplete information, and high stakes. If a model fails when answer choices are shuffled, how can we trust it with ambiguous symptoms or noisy imaging? Benchmarks don’t just evaluate models but shape how they’re trained and optimized [6]. When benchmarks reward test-taking shortcuts over genuine medical understanding, they create misleading signals of progress.
This paper exposes these fragilities through a series of targeted stress tests across six high-profile models and six widely used multimodal medical benchmarks. We show that state-of-the-art scores can hide medical brittleness and that model performance is often benchmark-specific, driven by exploiting artifacts rather than generalizable capability.
We then turn to the benchmarks themselves. Through structured clinician-guided analysis, we profile benchmark demands across reasoning complexity and visual dependency. The findings are striking: widely used benchmarks vary widely in what they actually test, yet are treated interchangeably in model evaluation. This masks key failure modes and risks misrepresenting real-world readiness.
Our findings call for a fundamental reevaluation of how we measure progress in health AI. Before these systems are used to support medical decision making, we must ensure they succeed for the right reasons, not merely because they can pass a test.
To that end, we apply a series of targeted stress tests that strip away spurious cues and shortcut opportunities. Even under these stricter conditions, performance improves across model generations—indicating real capability gains. But the gains are modest, and the brittleness persists. Further progress will require not just better models, but better ways to evaluate them.
2 Hidden Fragilities: Stress-Testing Models on Multimodal Medical benchmarks
High benchmark score performance may imply robustness, but actually conceals brittleness under distributional shift, missing data, or subtle perturbations. We designed a suite of stress tests that simulate uncertainty and reveal the behavioral shortcuts models may rely on. Rather than evaluating only correctness, these tests expose how models respond to degraded, incomplete, or adversarial input, conditions that characterize real-world use in healthcare environments.
We organize these stress tests into progressive levels of challenge, escalating from missing or misleading input modalities, to perturbations in structural format, and finally to qualitative audits of model reasoning.
Modality Robustness
Figure 2: Stress tests reveal overestimated modality robustness. Left (Stress Test 1). Accuracy drops when images are removed from two diagnostic benchmarks (NEJM and JAMA), revealing hidden differences in modality sensitivity across models. Sharp drops on NEJM—but minimal impact on JAMA—suggest inconsistent visual dependence across benchmarks and question types. Light segments show text-only accuracy; dark segments mark the drop from Image+Text. Right (Stress Test 2). Performance on 175 NEJM items requiring visual input. Most models exceed the 20% random baseline even without images, indicating reliance on shortcut cues such as priors, co-occurrence patterns, or memorized QA pairs. GPT-4o performs below chance due to refusal behavior.
Stress Test 1: Multimodal sensitivity
Objective: Assess how models respond to removal of image input on multimodal medical benchmarks.
[…]
Results: Removing images reveals strikingly different behavior across models that otherwise achieve similar benchmark scores. On NEJM, GPT-5 scored 80.89 with images and 67.56 without, a 13.33 pp drop. Similarly, Gemini 2.5 Pro dropped 14.94 pp (79.95 → 65.01), OpenAI-o3 fell by 13.86 (80.89 → 67.03), GPT-4o showed the largest drop (66.90 → 37.28, -29.62), and OpenAI-o4-mini showed a moderate drop (75.91 → 66.49, -9.42). DeepSeek-VL2 performed poorly in both modes (33.16 → 25.30).
On JAMA, absolute scores were higher and changes smaller: GPT-5 dropped 3.68 pp (86.59 → 82.91), Gemini 2.5 Pro dropped the most (84.84 → 74.93, -9.91), OpenAI-o3 dropped 2.10 (84.75 → 82.65), OpenAI-o4-mini dropped 2.10 (80.50 → 78.40), and GPT-4o dropped just 1.00 pp (69.90 → 68.90). The smallest deltas were seen in GPT-4o, suggesting either robustness to missing input or limited reliance on visual cues. Notably, DeepSeek-VL2 again underperformed (38.20 → 32.60). See Fig  2 and Extended Data Table 2 for details.
Interpretation: This test highlights two key issues. First, models with similar full-input scores diverge sharply when the image is removed, exposing hidden differences in modality sensitivity not reflected in benchmark leaderboards. Second, the benchmarks themselves vary in visual dependence: NEJM scores drop sharply without images, while JAMA scores remain relatively stable indicating that many items may be solvable from text alone. These benchmark-level inconsistencies motivate our item-level modality profiling in Section 3. To better understand whether score drops reflect robustness to missing input or reveal shortcut-based behavior, we next turn to questions where visual input is explicitly required.
Stress Test 2: Modality necessity
Objective: Test whether models can answer questions that require both text and image even when the image is missing, or whether their benchmark success reflects true multimodal ability.
Setup: We curated a 175-item subset from the NEJM benchmark, manually selecting multiple-choice questions that clinicians agreed require visual input to answer. We ensured that the selected items contained no textual cues, so the correct answer required both text and image. We then evaluated model performance under two conditions: text-only (image removed) and image+text. Models include GPT-5, Gemini-2.5 Pro, OpenAI-o3, OpenAI-4o-mini, GPT-4o, and DeepSeek-VL2. To ensure validity, we filtered out examples with content moderation triggers or ambiguous visual requirements.
Results: All models achieved high scores on image+text (e.g., GPT-5: 66.28%, Gemini-2.5 Pro: 67.42%, OpenAI-o3: 61.71%), confirming these items are solvable with full input. However, without the image, models still performed well above random chance (20%), despite the image being necessary. For example, GPT-5: 37.71%, OpenAI-4o-mini: 33.71%, Gemini-2.5 Pro: 37.14%, OpenAI-o3: 37.71%, suggesting models exploit non-visual cues or learn shortcut patterns from training data. GPT-4o, in contrast, scored just 3.4% due to frequent refusal to answer when the image was removed. While this deflates its accuracy, it may reflect more cautious behavior when critical visual input is missing, as opposed to relying on spurious textual cues. See Extended Data Table 3 for details.
Interpretation: This test isolates potential shortcut behavior by focusing only on visually required questions—cases where the text alone offers minimal diagnostic context (e.g., “What is the diagnosis?”). Yet models still perform well above chance without images, suggesting reliance on dataset artifacts—such as frequency priors, co-occurrence patterns, or memorized question–answer pairs—rather than genuine multimodal medical understanding. Unlike others, GPT-4o scored far below chance due to frequent abstention, which though penalized numerically, may better reflect appropriate uncertainty when essential visual context is missing (See Fig  7  8  9).
Importantly, this behavior does not imply model malpractice; many benchmark examples are publicly available online, making them susceptible to unintentional memorization during pretraining. More broadly, high benchmark performance may not reflect true multimodal understanding but instead arise from spurious cues or memorized associations, overstating model readiness for real-world medical use.
Shortcut Pattern Disruption
Stress Tests 1–2 revealed that LMMs often fall back on shortcuts when a modality is missing. We now examine the structure of these shortcuts more directly, introducing targeted input perturbations to break and analyze the superficial patterns underlying model predictions.
Figure 3: Stress tests uncover shortcut reliance under perturbation. Left (Stress Test 3). Option reordering causes accuracy to drop in the text-only setting (solid lines), exposing sensitivity to answer formatting. In contrast, accuracy is stable or improves under image+text input (dashed), suggesting vision helps override shallow textual heuristics. Middle (Stress Test 4). Accuracy declines as more distractors are replaced with irrelevant alternatives. All models approach random guessing when all distractors are substituted. Surprisingly, replacing a single distractor with an “Unknown” option boosts accuracy (+5–8pp), indicating models treat it as a weakened distractor—not an abstention signal. Right (Stress Test 5). Accuracy collapses when images are swapped to align with a distractor answer. Despite no textual changes, models fail to revise predictions—dropping over 30pp—highlighting brittle reliance on visual–label pairings rather than robust visual–textual reasoning. Line styles denote input modality: solid (text-only), dashed (image+text).
Stress Test 3: Format Perturbation
Objective: Evaluate whether models rely on answer presentation, such as fixed positions or option order, instead of question content when choosing answers.
Setup: We use the 175 manually curated NEJM items from Stress Test 2, where both text and image are required to answer correctly. The ground truth distribution is approximately uniform, minimizing class imbalance. We modify the multiple-choice options by randomly reordering them (without altering their meaning), to detect position-based bias. The correct answer remains unchanged in each case. We exclude DeepSeek-VL2 due to consistently meaningless outputs. Accuracy is measured under both text-only and image+text settings.
Results: Reordering distractors causes accuracy drops across all models in the text-only setting, suggesting reliance on surface-level patterns. For example, GPT-5 falls from 37.71% to 32.00%, Gemini 2.5 Pro from 37.14% to 33.14%, and OpenAI-o3 from 37.71% to 31.42%. In contrast, in the image+text setting, most models see a slight accuracy increase (e.g., GPT-5 from 66.28% to 70.85%), likely due to stochasticity or reduced distractor interference when vision is available. See Extended Data Table 4 for details.
Interpretation: This test reveals that models may rely on superficial answer formatting—particularly distractor position—when images are unavailable. In the text-only setting, shuffling the option order leads to consistent performance drops (e.g., GPT-5: 5.7-5.7 pp, Gemini 2.5 Pro: 4.0-4.0 pp, OpenAI-o3: 6.3-6.3 pp), suggesting shortcut behavior based on learned answer position patterns or co-occurrence biases, rather than medical understanding.
The ground truth label distribution is approximately uniform, and no answer content was altered—only distractor positions were randomized—ruling out class imbalance as an explanation. These results imply that models may have internalized fixed-format response from pretraining or prior benchmarks.
In the text+image setting, performance across most models remains stable or improves slightly. The three most recent models—GPT-5, OpenAI-o3, and Gemini 2.5 Pro—show modest gains (+2 to +4.6 pp), with GPT-5 improving the most. In contrast, GPT-4o and o4-mini remain largely unchanged. While these shifts are relatively small and may reflect stochastic variation, they suggest that visual inputs provide a more stable signal that helps mitigate brittle text-only heuristics. However, this does not imply that vision reliably overrides faulty reasoning—only that performance is less disrupted when additional modality cues are available.
Stress Test 4: Distractor Replacement
Objective: Test whether models rely on memorized question–option patterns, rather than genuine understanding, to identify answers.
Setup: We progressively replace 1–4 distractor options with irrelevant choices sampled from other items from the same dataset. The correct answer remains valid, but its label (e.g., “B”) may shift due to shuffling. We evaluate under both text-only and image+text conditions. We also include a variant where a single distractor is replaced with the token “Unknown”, effectively weakening the distractor. We reuse the 175 NEJM items from Stress Test 2.
Results: Under text-only input, model accuracy declines as more distractors are replaced with irrelevant options, approaching random guessing when all four are replaced. For example, GPT-5 accuracy drops from 37.7% (no replacement) to 20% (4 distractors replaced). Under image+text input, the trend is reversed: accuracy increases as more distractors are replaced. GPT-5 improves from 66.3% to 90.9% when all distractors are irrelevant.
A variant with a single distractor replaced by “Unknown” also improves performance slightly across models. For instance, GPT-5 improves by +5.25 percentage points. See Extended Data Table 5 for details.
Interpretation: This test examines how models behave when familiar distractors are weakened or removed. In the text-only setting, replacing distractors with irrelevant options leads to a consistent accuracy drop across all models. As more distractors are substituted, model confidence in correct guesses deteriorates toward chance—e.g., GPT-5 drops from 37.7% to 20.0% when all four distractors are replaced. This suggests that models rely heavily on learned question–option associations, frequency priors, or superficial patterns to make educated guesses in the absence of visual cues. While not unreasonable given minimal textual context (e.g., “What is the diagnosis?”), this behavior reveals a brittle dependency on dataset artifacts rather than genuine understanding.
Interestingly, replacing just one distractor with “Unknown” yields a consistent performance boost across all models (e.g., +5 to +8 percentage points), suggesting “Unknown” is not treated as a legitimate answer under incomplete input, but rather as a weakened distractor—effectively making the task easier by reducing ambiguity. Instead of abstaining or flagging uncertainty, frontier models become more confident in selecting the correct answer, further highlighting their elimination-based reasoning strategy.
In the image+text condition (full input), performance increases as distractors are replaced—e.g., GPT-5 rises from 66.3% to 90.9%, likely reflecting task simplification, as the substituted distractors are not visually aligned or grounded, and thus easier to rule out. Vision appears to become more helpful when spurious textual cues are disrupted.
Collectively, these observations suggest that apparent multimodal medical competence may arise from shallow elimination heuristics, with small perturbations exposing fundamental fragilities in model behavior.
Stress Test 5: Visual Substitution
Objective: Evaluate whether models truly integrate visual and textual information, or instead rely on shallow visual–answer associations learned during training.
Setup: We manually curated 40 NEJM examples where the correct answer is tightly linked to specific visual features. For each item, we substitute the original image with one that aligns with a distractor choice, while keeping the question and answer options unchanged. This creates a new “counterfactual” version where the distractor becomes the intended correct answer. A model with genuine visual–textual understanding should revise its prediction in response to the changed image. We measure performance under this perturbed image+text setting and compare against the original.
Results: Model performance drops sharply in this altered setting—despite the minimal input change. For example: GPT-5 drops from 83.3% to 51.7% ($\Delta$ -31.6 pp), and Gemini-2.5 Pro falls from 80.8% to 47.5% ($\Delta$ -33.3 pp). OpenAI-o3 and o4-mini show similar declines (e.g., o4-mini: -34.2 pp). GPT-4o is an exception, showing minimal change.
These sharp declines suggest that many models rely on learned visual–answer pairings rather than interpreting visual evidence in context. See Extended Data Table 6 for details, with illustrative test cases provided in Appendix B.
Interpretation: This test introduces a targeted perturbation designed to minimally alter the question while shifting the intended answer via the image. The substantial performance drop across models highlights a brittle reliance on visual–label associations—rather than robust visual–textual integration.
We acknowledge that this perturbation may also change task difficulty, as the new images vary in visual salience or diagnostic ambiguity.
Nonetheless, the consistent accuracy collapse—despite no change in textual input—suggests that current benchmarks may reward shortcut behavior over genuine visual reasoning. The fact that performance collapses despite unchanged text reveals that high benchmark scores may reflect brittle visual–answer shortcuts, rather than reliable visual reasoning or robust multimodal understanding.

### v3

# The Illusion of Readiness in Health AI

@@ Abstract
Large language models have demonstrated remarkable performance in a wide range of medical benchmarks. Yet underneath the seemingly promising results lie salient growth areas, especially in cutting-edge frontiers such as multimodal reasoning. In this paper, we introduce a series of adversarial stress tests to systematically assess the robustness of flagship models and medical benchmarks. Our study reveals prevalent brittleness in the presence of simple adversarial transformations: leading systems can guess the right answer even with key inputs removed, yet may get confused by the slightest prompt alterations, while fabricating convincing yet flawed reasoning traces. Using clinician-guided rubrics, we demonstrate that popular medical benchmarks vary widely in what they truly measure. Our study reveals significant competency gaps of frontier AI in attaining real-world readiness for health applications. If we want AI to earn trust in healthcare, we must demand more than leaderboard wins and must hold AI systems accountable to ensure robustness, sound reasoning, and alignment with real medical demands.
Abstract
submission/sections/abstract
Introduction
Large language models (LLMs) have demonstrated impressive health AI capabilities such as passing medical exams and acing diagnostic benchmarks [1, 2, 3, 4, 5, 6]. But these promising results belie the underlying fragility and competency gaps [7, 8]. In this paper, we evaluate the readiness of frontier AI models in performing health-related multimodal reasoning. While prior work abounds in studying LLM pitfalls [9, 10, 11], they predominantly focus on the text modality, whereas multimodal reasoning is relatively underexplored. As multimodal generative AI has emerged as the next frontier for biomedicine [12], it is imperative to bridge this gap.
Inspired by the study of adversarial learning that unveiled the underlying fragility of first-gen deep learning [13, 14], we introduce six stress tests on the multimodal reasoning capabilities of health AI models by systematically perturbing test instances in popular medical benchmarks. Our evaluation uncovers surprising failure modes of LLMs. They can produce the right answer even when not supposed to (e.g., when the input image was removed). On the other hand, their performance may collapse even with the slightest distraction (e.g., when the answer choices were shuffled). Finally and perhaps most concerning for interpretability and accountability, when pressed for justification, they fabricate convincing yet flawed reasoning traces. Collectively, these stress tests reveal a rather different picture underneath the seemingly consistent progress LLMs have appeared to attained in multimodal medical benchmarks.
These findings inspire us to further benchmark the benchmarks themselves. Specifically, we introduce clinician-guided rubrics to tease apart the various aspects a test instance is measuring, from image perception to reasoning. This reveals interesting diversity among popular multimodal medical benchmarks, which are often measuring rather different things but are treated interchangeably.
Our results call for a fundamental re-evaluation of how to measure progress in health AI. To this end, we compile a harm ontology for multimodal reasoning in health AI. There is great potential for LLMs and generative AI to empower clinical practitioners and patients alike, but before these systems are deployed to end-to-end clinical applications, we must ensure they are passing the right tests that can truly assess a system’s readiness in real-world scenarios.
Results
Stress testing exposes robustness gaps in LLMs
LLMs achieve high scores on multimodal medical benchmarks, often interpreted as evidence of robust clinical competence. However, these aggregate scores may conceal brittleness under incomplete, perturbed, or ambiguous inputs, conditions that commonly arise in real-world healthcare settings. To assess model behavior beyond correctness alone, we developed a suite of six stress tests probing how models respond to degraded input modalities, structural variations, distractors, and reasoning demands.
Rather than evaluating accuracy in isolation, these tests also measure model abstention behavior, reasoning fidelity, and stability under perturbation. This allows us to identify not just performance drops, but how and why model outputs change, exposing brittle dependencies, shortcut behaviors, and inconsistency patterns that standard benchmarks may overlook.
The stress tests are organized to escalate in challenge: from missing or misleading inputs (Stress Tests 1-2), to format-based perturbations (Tests 3-4), to altered visual evidence (Test 5), and finally to direct audits of model-generated reasoning (Test 6). Figure 1 summarizes the framework and illustrates representative model responses under stress. Accuracy is defined as selection of the original ground-truth answer prior to perturbation.
Figure 1: Stress tests expose hidden fragility and salient robustness gaps of large language models (LLMs) in multimodal health AI tasks. a, Strong performance in standard benchmarks belies underlying fragilities in health AI by LLMs. b, Seemingly strong performance in standard benchmarks (green) contrast with robustness gaps (red), as a suite of stress tests showed that LLMs often produce incorrect answers confidently and get distracted easily. c, Robustness scores reveal salient growth areas in LLMs, including the more recent reasoning models. d, A representative example showing how GPT-5 guessed the correct answer even without the input image and justified it with faulty reasoning, illustrating why it’s important to go beyond standard benchmark scores and introduce stress tests to assess the real understanding of LLMs in health AI tasks.
Figure 2: Stress tests reveal modality sensitivity and shortcut behavior in large language models. T1: Modality sensitivity on multimodal diagnostic benchmarks. a, Accuracy on NEJM and JAMA benchmarks with full (image+text) and text-only inputs. Bars show full-input accuracy; lighter segments indicate performance drop after image removal. Text-only performance remains high, suggesting that benchmark scores may not reflect true visual understanding. T2: Behavior without images on the NEJM Visual-Required Subset (NEJM-VS). This 175-question subset was independently selected by board-certified clinicians to require visual interpretation. b, Despite lacking images, models achieve well above the 20% chance level, indicating potential reliance on non-visual cues or memorized associations. Accuracy includes abstentions counted as incorrect. Notably, when excluding abstentions, GPT-4o achieves $\sim$40 % accuracy, on par with other models, suggesting similar levels of shortcut success. Its much higher abstention rate may reflect a more cautious or instruction-sensitive response pattern. c, Answer-type breakdown reveals models often provide confident but incorrect responses or shortcut-derived correct guesses. Ideally, models should abstain or fall back to random guessing in the absence of required visual input.
Modality sensitivity and input omission
Stress Test 1: Input removal. We evaluated model performance under full input (image + text) and text-only conditions on two multimodal medical benchmarks: NEJM [15] and JAMA [16]. Each question pairs a short clinical vignette with one or more diagnostic images (radiology, dermatology, or pathology).
Across six models, removing the image led to consistent performance declines on NEJM, with paired accuracy differences ranging from -9.29 to -27.05 percentage points (pp). For example, GPT-5 [3] scored 80.89% (95% CI 78.06-83.71) with full input and 67.56% (64.20-70.93) without ($\Delta=-13.33$ pp), while GPT-4o [17] dropped from 64.33% (60.83-67.83) to 37.28% (33.78-40.78) ($\Delta=-27.05$ pp). On JAMA, declines were smaller overall: GPT-5 decreased 3.68 pp (86.59% → 82.91%), and GPT-4o decreased 1.05 pp (69.94% → 68.89%). See Fig. 2 and Extended Data Table 3.
Stress Test 2: Visual-necessity subset. To test whether models can still answer questions that explicitly require image information, we curated a 175-item subset of the NEJM benchmark, referred to as NEJM Visual-required Subset (NEJM-VS), comprising cases selected based on clinical criteria indicating minimal textual cues and a high dependency on visual features.
With full input (image + text), models achieved moderate to high accuracy-GPT-5 66.3% (95% CI 59.4-73.1), Gemini-2.5 Pro [18] 67.4% (60.0-74.3), and OpenAI-o3 [19] 61.7% (54.9-68.6)-confirming that these items were solvable when the image was available. In the text-only condition, accuracy remained well above the 20% random baseline: GPT-5 37.7% (30.3-45.1), Gemini 37.1% (30.3-44.6), OpenAI-o3 37.7% (30.3-45.1), and OpenAI-o4-mini [19] 33.7% (26.9-40.6). Despite the image being essential, these models maintained performance nearly twice the chance level, indicating reliance on non-visual cues or memorized associations. In contrast, GPT-4o achieved only 3.4% (1.1-6.3) accuracy, reflecting a refusal rate of roughly 97% when images were removed. Accuracy includes abstentions counted as incorrect. When abstentions are excluded, GPT-4o’s accuracy rises to $\sim$40%, comparable to other models, suggesting that its lower overall score reflects conservative uncertainty handling, rather than diminished shortcut behavior. See Extended Data Table 4 for complete results.
Interpretation. Most models scored well above random chance level despite the absence of images, which may be the result of reliance on non-visual cues (i.e. disease prevalence) or memorized associations (shortcut learning). Ideally, model should recognize when critical input is missing and warn that predictions may be unreliable. GPT-4o’s abstention behavior, though numerically penalized, demonstrates more appropriate uncertainty handling. Together, these results expose wide variation in modality sensitivity among models that otherwise achieve similar headline benchmark scores.
Figure 3: Stress tests reveal shortcut reliance under input perturbations. Each column corresponds to a stress test (T3–T5); the top row (a–d) shows accuracy with image+text input, and the bottom row (e–g) shows accuracy with text-only input. Accuracy is defined as selection of the original ground-truth answer prior to perturbation. Error bars indicate 95% confidence intervals. T3: Format perturbation (a, e). Randomizing answer order weakens positional cues. In the text-only condition (e), performance declines, suggesting reliance on answer formatting. In contrast, accuracy with full input (a) is preserved or improved, likely due to increased reliance on visual grounding when shortcuts are removed. T4: Distractor replacement. T4a (b, f): One distractor is replaced with “Unknown.” Most models improve accuracy even in the text-only setting (f), where “Unknown” should serve as a fallback when visual information is missing. This indicates models use “Unknown” as a semantic cue rather than abstaining, a failure of robust fallback behavior. T4b (c, g): One to four distractors are iteratively replaced with answer choices drawn from unrelated questions. Accuracy in the text-only condition (g) drops gradually toward the 20% chance baseline, consistent with progressive shortcut removal. With image+text input (c), accuracy increases across models, reflecting improved disambiguation from visual features. T5: Visual substitution (d). The original image is replaced with a distractor-matching image. Accuracy drops sharply across models, despite unchanged text input, demonstrating reliance on visual features and vulnerability to visual perturbation.
Shortcut reliance under perturbation
Stress Test 3: Format sensitivity. To assess whether models rely on superficial formatting rather than question content, we randomized the order of answer choices without altering the question text or correct answer. On the NEJM Visual-Required Subset (NEJM-VS), accuracy under text-only input declined across all models (e.g., GPT-5: –5.7 pp, Gemini: –4.0 pp, o3: –6.3 pp), indicating sensitivity to positional biases. By contrast, performance under image+text input remained stable or improved slightly (typically +2–5 pp), suggesting that visual information compensates for disrupted textual cues. Full confidence intervals are shown in Extended Data Table 5.
Stress Test 4: Distractor manipulation. We next tested whether models rely on shortcut patterns involving familiar distractors. For each question, 1–4 incorrect choices were replaced with unrelated options, and in one variant, a distractor was replaced with “Unknown.”, serving as a functional “unknown” option.
Under text-only conditions, accuracy declined steadily toward the 20% chance level as more distractors were replaced (e.g., GPT-5: –17.7 pp, Gemini: –20.0 pp), consistent with shortcut removal. Conversely, in the image+text condition, performance improved markedly (up to +25 pp), as irrelevant distractors made the correct answer more visually salient. When a single distractor was replaced with “Unknown,” accuracy increased in both modalities-most notably under text-only input-suggesting that models treat “Unknown” as an easily eliminated option rather than a fallback when uncertain. Detailed results are provided in Extended Data Table 6.
Interpretation. These perturbation-based tests reveal that models often depend on superficial or distributional patterns when visual information is unavailable. Accuracy degrades when positional or semantic cues are disrupted, even though the question and correct answer remain unchanged. In contrast, access to visual input stabilizes performance, enabling models to bypass these fragile shortcuts. However, improvements with “Unknown” and rapid recovery from distractor replacement suggest reliance on elimination strategies rather than true visual reasoning. This highlights a critical risk: high accuracy in benchmark settings may overstate a model’s multimodal robustness, masking shortcut reliance under minor perturbations.
Multimodal grounding failures
Stress Test 5: Visual substitution. To assess whether models engage in adaptive multimodal reasoning, we designed a controlled substitution task targeting visual grounding. We curated 40 NEJM questions whose correct diagnoses depended critically on image interpretation, then replaced each original image with a clinically plausible alternative corresponding to a distractor option. The vignette and answer choices remained unchanged. All substituted images were independently verified by clinicians to support the newly assigned ground-truth answer. Models were evaluated under the image+text condition only.
Despite minimal textual change, performance declined markedly across most models. GPT-5 accuracy dropped from 83.3% to 51.7% (–31.6 pp), Gemini-2.5 Pro from 80.8% to 47.5% (–33.3 pp), o4-mini by –34.2 pp, and o3 by –24.2 pp. GPT-4o was the only model to maintain similar performance (63.8% vs. 66.7%).
Extended results are presented in Extended Data Table 7, with qualitative examples in Appendix Supplementary Note 2 | More stress test results.
Interpretation. This stress test isolates visual grounding by altering only the diagnostic image while preserving the textual input. A reliable LLM should revise its answer when the visual evidence changes. However, most models failed to adjust properly, suggesting reliance on static image–answer pairings or incomplete visual-text integration. The large performance drops (30–35 pp) indicate that standard benchmark accuracy may overstate multimodal reasoning capacity.
Residual performance in the perturbed setting (e.g., GPT-5 at 51.7%) may reflect partial recognition of generic visual features, model overconfidence, or memorization of common image–question associations during pretraining. Although all substituted images were drawn from publicly available datasets and validated for clinical correctness, we cannot fully exclude overlap with model pretraining corpora.
These findings highlight a critical limitation: while LLMs often detect that the image is relevant, they fail to dynamically reinterpret its meaning in context. Such brittle visual grounding undermines diagnostic trustworthiness and reveals a key gap between benchmark success and robust clinical decision-making.